A two storeyed building has the height of its lower storey 12 metre and that of upper storey 13 metre. Find at what distance the two storeys subtend equal angles to an observer's eye at height 5 metre from the ground.
1
Expert's answer
2012-11-29T08:52:57-0500
Solution:
Today's price equal:
P = D*(1-g)/(k-g) = D*(1-g)/(rfr+beta(rm - rfr)-g)= $1.25*1/(0.03+1.2(0.055))=
$1.25/0.096 = $13.02

Answer: $13.02